OpenStudy (anonymous):

how do you solve t=5m+5v

OpenStudy (johnweldon1993):

depends on what you want to solve for....t, m, or v?

OpenStudy (anonymous):

For M

OpenStudy (johnweldon1993):

so then you would just rearrange the order to make m = something so we have t = 5m + 5y we want 5m alone for now...so subtract 5y from both sides t - 5y = 5m now we want m totally by itself because it didnt ask to solve for 5m lol so divide everything by 5 t/5 - 5/5 = 5m/5 t/5 - 1 = m so m would = t/5 - 1

OpenStudy (anonymous):

thanks

OpenStudy (johnweldon1993):

no problem!